Two point charges are fixed on the y axis: a negative point charge q1 = -30 μC at y1 = +0.17 m and a positive point charge q2 at y2 = +0.30 m. A third point charge q = +8.8 μC is fixed at the origin. The net electrostatic force exerted on the charge q by the other two charges has a magnitude of 24 N and points in the +y direction. Determine the magnitude of q2.
Answer #1


the force F1 between q1 and q is F1 = k*q1*q/y1^2 = (9*10^9*30*10^-6*8.8*10^-6)/0.17^2 = 82.2 N along +Y-axis

the Force between q2 and q is F2 = k*q2*q/y2^2 = (9*10^9*q2*8.8*10^-6)/(0.3^2) = 8.8*10^5*q2

this should be positive charge to get force in -Y direction

Net force F = F1+F2 = 24 N

82.2 - (8.8*10^5*q2) = 24

q2 = 66.1*10^-6 C = 66.1 uC